OpenStudy (anonymous):

A rectangular prism has a volume of 260 in.³ Its width is 5 in. and its length is 8 in. The height of the prism is _____ in.

jimthompson5910 (jim_thompson5910):

V = LWH 260 = 8*5*H 260 = 40H 260/40 = H 6.5 = H H = 6.5 So the height is 6.5 inches